OpenStudy (anonymous):

I am standing at point A on one side of a river and wish to measure the distance across a river to a house. H on the other side. I walk 50 feet along the riverbank to point B and measure the angle ABH to be 18°. Find the distance across the river(assume a right triangle).

OpenStudy (anonymous):



OpenStudy (anonymous):

IS THIS RIGHT

OpenStudy (anonymous):

Its quite simple. Draw the right angled triangle. You'll find that tan(18)= 50/x where x is the distance you need.

OpenStudy (anonymous):

The 18 degrees should be at H :)

OpenStudy (anonymous):

You got the answer?

OpenStudy (anonymous):

first cos 18=50/hyp find hyp and apply pathagoroous to find x or sin 18=x/hyp